OpenStudy (anonymous):

The graphs of the function F (left, in blue) and G (right, in red) are below. Let P(x)=F(x)G(x) and Q(x)=F(x)/G(x). Answer the following questions. P'(1) Q'(1) P'(6) Q'(6)

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

i have this F'(1)=0 F(1)=1 G'(1)=1 G(1)=2 F'(6)=1/4 F(6)=3 G'(6)=3 G(6)=3

OpenStudy (anonymous):

That all looks right to me

OpenStudy (anonymous):

is it saying you're wrong?

OpenStudy (anonymous):

but i have to figure out P'(1) etc. i cant get those

OpenStudy (anonymous):

So what do you think the first one, P'(x) would be?

OpenStudy (anonymous):

I got P'(1)=1

OpenStudy (anonymous):

the formula! numbers should always be an afterthought :) How did you get 1?

OpenStudy (anonymous):

fg'+gf'

OpenStudy (anonymous):

yup. F'(1) = 0, so it's just F(1)G'(1) = 1*3

OpenStudy (anonymous):

g'(1)=1 not three

OpenStudy (anonymous):

you're totally right. I misread. Then I'd think that's right. It says no?

OpenStudy (anonymous):

all of them have to be right in order to get the point. so at least one of the others is wrong

OpenStudy (anonymous):

Q'(1)=\[(gf'-fg')/(g^2)\]

OpenStudy (anonymous):

i got -1/2

OpenStudy (anonymous):

g^2 on the bottom

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-1*1/(2^2) g(1) = 2

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

never mind its -1/4

OpenStudy (anonymous):

what about the other two P'(6) and Q'(6)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What'd you get for em?

OpenStudy (anonymous):

39/4, -11/12

OpenStudy (anonymous):

I got 3/2 and -7/8 ...

OpenStudy (anonymous):

but I'm way wrong, one sec. keep mixing things up

OpenStudy (anonymous):

7/4 for the first 3/4 + 1

OpenStudy (anonymous):

shouldnt it be 9+3/4

OpenStudy (anonymous):

F'(1)=0 F(1)=1 G'(1)=1 G(1)=2 F'(6)=1/4 F(6)=3 G'(6)=3 G(6)=3

OpenStudy (anonymous):

yeah, you're right. I'm terrible at this right now :P

OpenStudy (anonymous):

so P'(6) is 39/4

OpenStudy (anonymous):

the second one I got 1/12 - 1 but I'm wholly unreliable so far :P

OpenStudy (anonymous):

yeah, 39/4 should be right

OpenStudy (anonymous):

what about Q'(6)

OpenStudy (anonymous):

I still got -11/12 just like you had

OpenStudy (anonymous):

somethings is still wrong

OpenStudy (anonymous):

G is red and F is blue, right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

really wish that it would tell you if each part was right.... hmmmm...

OpenStudy (anonymous):

I can find nothing wrong...

OpenStudy (anonymous):

I'm sorry :(

OpenStudy (anonymous):

oh well we tried. i go to a study session on mondays i will get it done their. thanks

OpenStudy (anonymous):

Good luck. I'll keep staring at this for a sec to see if anything comes of it :/

OpenStudy (anonymous):

Do you have to put inputs in an online form sort of thing?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

did you try entering them as decimals? haha... hmm..

OpenStudy (anonymous):

that doesnt matter

OpenStudy (anonymous):

they worked on the other problem, so I figured that wasn't the problem